OpenStudy (anonymous):

Assume that oil spilled from a ruptured tanker spreads in a circular pattern whose radius increases at a constant rate of 2ft/s. How fast is the area of the spill increasing when the radius of the spill is 60ft? formula=A(PI)r^2

OpenStudy (anonymous):

A = pi r^2 -> A' = 2 pi *r * r' = 2 pi*60* 2 = 754 ft^2/sec